# Long-Lived Photon Blockade with Weak Optical Nonlinearity

### v2

@@ Abstract
Antibunching time window—In Fig. 2, the green curve shows the time-delayed second-order correlation function $g^{(2)}(\tau)$ for the above four-cavity system, at one of its LLPB points (at $\gamma=1$). The photon antibunching time window is approximately $8/\gamma$. This plot is obtained from WFMC simulations, but the analytical theory gives almost identical results. The analytic approximation for this result can be derived for $J^{\prime}\ll J$ by approximating the eigenvalues and eigenvectors of the coupling Hamiltonian to second order in $J^{\prime}$. The full result is given in Supplemental Materials [55]. If we look into the short time scale $\tau\sim 1/\gamma$ and expand $g^{(2)}_{22}(\tau)$ as polynomial of $\gamma\tau$, assuming negligible $\alpha$, we find
$g_{22}^{(2)}(\tau)$ $\approx(\gamma\tau)^{4}/64.$ (23)
This small $\tau$ expansion is order-4 and fundamentally distinct from the quadratic conventional photon blockade case $g^{(2)}(\tau)\approx(\gamma\tau)^{2}/4$, by Eq. (14).
For comparison, the blue curve in Fig. 2 shows $g^{(2)}(\tau)$ for a comparable two-cavity UPB setup [30]. Here we take $\alpha=0.001227$, the same as in the four-cavity LLPB case. For the same decay rate $\gamma=1$, two-cavity UPB is achieved for $J\approx 17.67$, and the strong coupling results in rapid oscillations of $g^{(2)}(\tau)$. The central antibunching window, estimated as the window over which $g^{(2)}(\tau)<0.5$, is around 90 times smaller than in the LLPB case.
As a further comparison, the orange dashes in Fig. 2 show $g^{(2)}(\tau)$ for a one-cavity setup with $\gamma=1$ and $\alpha=10$. (Note that $g^{(2)}(0)$ does not reach precisely zero under conventional photon blockade; we simply choose $\alpha$ so that it is visibly suppressed.) The $g^{(2)}(\tau)$ curves for both LLPB and conventional photon blockade have no oscillations, unlike the two-cavity UPB case. Remarkably, the LLPB curve has a significantly flatter bottom near $\tau=0$ than the curve for conventional photon blockade, and the time window is larger by a factor of around 1.68. This can be understood from Eq. (23); the expansion leading term is of order-4 in $\gamma\tau$, approximately $(\gamma\tau)^{4}/64$, compared to the quadratic term $\approx(\gamma\tau)^{2}/4$ for the single-cavity case. This characteristic results in a more gradual change in the dip, which may be advantageous for pulsed device operations [38, 41].
@@
Conclusion—We have re-analyzed the origins of unconventional photon blockade (UPB), in which photon antibunching is achieved by quantum interference in weakly nonlinear coupled cavities [30]. A major problem with UPB has hitherto been the rapid oscillation in $g^{(2)}(\tau)$ due to strong inter-cavity coupling [38, 41]. We show that this can be overcome by designing a coupled-cavity system so that the linear single-photon Green’s function exhibits a zero (a 1-photon dark state) at a large cavity loss rate. This leads to a “long-lived photon blockade” (LLPB) scenario whereby $g^{(2)}(\tau)\propto\tau^{4}$ and the time window exceeds that of conventional photon blockade under strong nonlinearity. It is remarkable that shortening the cavity lifetime can effectively lengthen the antibunching time window, which can be regarded as a fresh example of non-Hermiticity (in the form of dissipation) producing unexpected physical behaviors [65, 66]. Our proposal does not need one-way inter-cavity couplings or other exotic requirements [53, 41], and should be compatible with quantum optics experiments using setups like the one shown in Fig. 1(c) [55].
@@
Our theoretical framework offers ample scope for further generalization and optimization. For instance, in the four-cavity configuration of Fig. 1(b), the couplings are described by just three parameters for simplicity, but they can instead be individually optimized to achieve an even larger $g^{(2)}(\tau)$ time window. Coherently driving multiple cavities also offers additional possibilities for achieving LLPB with other cavity configurations. Possible implementation challenges may include the need for fine-tuning and ensuring the stability of the inter-cavity couplings and other system parameters, which is a problem shared with the original UPB setup and all other related schemes to date.
@@
where $G_{kd}\equiv\langle k|G|d\rangle$ is the $k,d$ element of the single-photon Green’s function and
@@
$\langle i|\psi_{j}^{(1)}(\tau)\rangle=\langle j|\bar{\psi}^{(1)}\rangle\langle i|\bar{\psi}^{(1)}\rangle-2\alpha F_{d}^{2}\sum_{k}G_{kd}^{2}\langle i,j|G^{(2)}(\tau)|k,k\rangle,\$ (S18)
@@
$G^{(2)}(\tau)\equiv\sum_{mn}\frac{|\varphi_{mn}\rangle\langle\varphi_{mn}|}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i(z+\epsilon_{n})\tau}.$ (S19)
@@
$=\left|1-2\alpha\sum_{k}\frac{G_{kd}^{2}}{G_{id}G_{jd}}\langle i,j|G^{(2)}(\tau)|k,k\rangle\right|^{2}.$ (S21)
@@
First, we will develop some intuition about how photon antibunching arises in this system. In the Dyson series expansion for $G_{id}(z)$, each term in the series can be interpreted as a photon trajectory starting at the driven cavity $d$ and ending at the output cavity $i$ [60]. Along a trajectory, each cavity contributes $1/z$ to the Green’s function. For $G_{21}$, the Dyson series up to order $1/z^{4}$ has five terms, corresponding to the following trajectories:
@@
S6 Photonics Realization

### v3

@@ Abstract
Antibunching time window—In Fig. 2, we plot the time-delayed correlation function $g^{(2)}(\tau)$ versus $\tau$, for the above four-cavity system at the $\gamma=1$ LLPB point (green curve). The photon antibunching time window, which we characterize by $g^{(2)}(\tau)<0.5$, is $\delta\tau\approx 8/\gamma$. This plot is obtained from WFMC simulations, but the theory based on Eqs. (9)–(11) gives almost identical results [55]. For comparison, the blue curve in Fig. 2 shows $g^{(2)}(\tau)$ for an analogous two-cavity UPB setup [30]. With the same nonlinearity strength $\alpha$ and decay rate $\gamma$ used in the four-cavity LLPB result, the two-cavity system achieves UPB for $J\approx 17.67$. This results in rapid oscillations of $g^{(2)}(\tau)$, with an antibunching window (estimated in the same way) around 90 times smaller than under LLPB.
As a further comparison, the orange dashes in Fig. 2 show $g^{(2)}(\tau)$ for a one-cavity setup with $\gamma=1$ and $\alpha=10$. (Note that $g^{(2)}(0)$ does not reach precisely zero under conventional photon blockade; we simply choose $\alpha$ so that it is visibly suppressed.)
Returning to the LLPB result, we also note that $g^{(2)}_{22}(\tau)$ scales very slowly for small $\tau$. To quantify this, we expand it as a polynomial in $\gamma\tau$ for $\gamma\tau\ll 1$. We obtain
$g_{22}^{(2)}(\tau)$ $\approx(\gamma\tau)^{4}/64,$ (19)
which is indeed much slower than the UPB and conventional photon blockade results, both of which scale as $\tau^{2}$ [55]. This “flat bottom” feature of LLPB may be particularly advantageous for pulsed device operations [38, 41].
@@
Conclusion—We have re-analyzed the origins of unconventional photon blockade (UPB), in which photon antibunching is achieved by quantum interference in weakly nonlinear coupled cavities [30]. A major problem with UPB has hitherto been the rapid oscillation in $g^{(2)}(\tau)$ due to strong inter-cavity coupling [38, 41]. We show that this can be overcome by designing a coupled-cavity system so that the linear single-particle Green’s function exhibits a zero (an SPDS) at a large cavity loss rate. This leads to a “long-lived photon blockade” (LLPB) scenario whereby $g^{(2)}(\tau)\propto\tau^{4}$ and the time window exceeds that of conventional photon blockade under strong nonlinearity. It is remarkable that shortening the cavity lifetime can effectively lengthen the antibunching time window, which can be regarded as a fresh example of non-Hermiticity (in the form of dissipation) producing unexpected physical effects [65, 66]. Our proposal does not need one-way inter-cavity couplings or other exotic requirements [53, 41], and should be compatible with quantum optics experiments using setups like the one shown in Fig. 1(c) [55].
@@
This theoretical framework offers ample scope for further generalization and optimization. For instance, in the four-cavity configuration of Fig. 1(b), the couplings are described by just three parameters for simplicity, but they can instead be individually optimized to achieve an even larger $g^{(2)}(\tau)$ time window. Coherently driving multiple cavities also offers additional possibilities for achieving LLPB with other cavity configurations. Possible implementation challenges may include the need for fine-tuning and ensuring the stability of the inter-cavity couplings and other system parameters, which is a problem shared with the original UPB setup and all other related schemes to date.
@@
where $G_{kd}\equiv\langle k|G|d\rangle$ is the $k,d$ element of the single-particle Green’s function and
@@
$\langle i|\psi_{j}^{(1)}(\tau)\rangle=\langle j|\bar{\psi}^{(1)}\rangle\langle i|\bar{\psi}^{(1)}\rangle-2\alpha F_{d}^{2}e^{-iz\tau}\sum_{k}G_{kd}^{2}\langle i,j|G^{(2)}(\tau)|k,k\rangle,\$ (S18)
@@
$G^{(2)}(\tau)\equiv\sum_{mn}\frac{|\varphi_{mn}\rangle\langle\varphi_{mn}|}{2z+\epsilon_{m}+\epsilon_{n}}e^{-i\epsilon_{n}\tau}.$ (S19)
@@
$=\left|1-2\alpha e^{-iz\tau}\sum_{k}\frac{G_{kd}^{2}}{G_{id}G_{jd}}\langle i,j|G^{(2)}(\tau)|k,k\rangle\right|^{2}.$ (S21)
@@
First, we will develop some intuition about how photon antibunching arises in this system. In the Dyson series expansion for $G_{id}$ as a function of $z$, each term in the series can be interpreted as a photon trajectory starting at the driven cavity $d$ and ending at the output cavity $i$ [56]. Along a trajectory, each cavity contributes $1/z$ to the Green’s function. For $G_{21}$, the Dyson series up to order $1/z^{4}$ has five terms, corresponding to the following trajectories:
@@
S6 WFMC simulation details
Wave function Monte Carlo (WFMC) simulations [61, 62, 63, 64] were performed using the QuantumOptics.jl toolbox [67], operating in the tensor product of truncated Fock spaces across lattice sites. Due to weak driving and uneven photon occupations, we adopted the following strategies to ensure accuracy and efficiency:
  • To avoid undersampling rare but potentially important jump events, we modified the standard master equation. For jump operators $L_{j}=\sqrt{\gamma}a_{j}$, we applied the transformation:
    $\mathcal{H}$ $\rightarrow\mathcal{H}^{\prime}=\mathcal{H}+\frac{\beta}{2i}\sum_{j}(L_{j}-L_{j}^{\dagger})$ (S59)
    $L_{j}$ $\rightarrow L^{\prime}_{j}=L_{j}+\beta\mathcal{I}$ (S60)
    with $\beta=0.1$. This increases jump rates while preserving the correct ensemble evolution and is known as choosing a different unraveling [54].
  • Simulations were performed with tolerances $\texttt{RelTol}=\texttt{AbsTol}=10^{-21}$.
  • To capture steady-state behavior, each trajectory was evolved for a relaxation time $T_{\text{relax}}$, followed by a recording period $T_{\text{record}}$. We used $T_{\text{relax}}=100$, $T_{\text{record}}=1000$, and the number of trajectories $N_{\text{traj}}=10$.
  • Fock cutoffs $N_{\text{Fock}}^{(j)}$ were chosen per site to balance accuracy and efficiency. For the driven cavity, we used $N_{\text{Fock}}^{(1)}=16$, while for the other cavities, we used $N_{\text{Fock}}^{(j\neq 1)}=8$.
Convergence was verified in all cases.
S7 Photonics Realization